The wave length of a travelling wave is 5m at a frequency of 12Hz and has an amplitude of 15m, write down the equation of wave

1
Expert's answer
2022-02-11T08:48:47-0500

Given:

"\lambda=5\:\rm m"

"f=12\:\rm Hz"

"A=15\:\rm m"


 The equation of wave is given by

"y(x,t)=A\sin\left(\frac{2\pi}{\lambda}x-2\pi f t\right)"

"y(x,t)=(15\:{\rm m})\sin\left(\frac{2\pi}{5}x-2\pi *12 t\right)"

"y(x,t)=(15\:{\rm m})\sin\left(0.4\pi x-24\pi t\right)"